# CACL: Community-Aware Heterogeneous Graph Contrastive Learning for Social Media Bot Detection

### v1

@@ 4 Experimental Setups
We set $\lambda$ to 0.9, $\tau$ to 0.07, learning rate to 0.0001, layer number of all GNNs models to 2, and dropout rate to 0.5. In each epoch, we iteratively select 2000 adjacent users for Twibot-22 datasets and 1000 for other datasets. We conduct all experiments on 2 NVIDIA RTX A5000 GPUs.
@@ 5.3 Community-aware Study
5.4 Ablation Study
@@ 6 Conclusion
In this paper, we mainly focus on constructing a community-aware graph contrastive learning framework dubbed CACL. Despite our best efforts, this paper may still have some limitations. We evaluate our method in three benchmarks, including the most comprehensive dataset Twibot-22 in Twitter bot detection area. However, Hays et al. (2023) still point out that current datasets still have certain limitations in collecting and labeling that affect the evaluation of methods. If we find several communities in real world and record the changes of them during the training process, we can simply reveal the effectiveness of our proposed framework. However, due to our framework is dynamic updated, community structure is also dynamic which brings problem to recording.
@@
Our research in this paper follows ethical guidelines and principles. All experiments and data usage are conducted in accordance with the relevant institutional and national guidelines and regulations. Informed consent was obtained from all participants involved in data collection or annotation tasks. Despite the datasets we use contain a large amount of private information existing in social media, we ensure the privacy and confidentiality of any personal or sensitive information used and obtained during the study. Our work aims to contribute positively to online environment, while respecting the dignity, rights, and well-being of individuals and communities. We acknowledge and address any potential biases or ethical considerations in our research.

### v3

@@ 4 Experimental Setups
We set $\lambda$ to 0.9, $\tau$ to 0.07, learning rate to 0.0001, layer number of all GNNs models to 2, and dropout rate to 0.5. In each epoch, we iteratively select 2000 adjacent users for Twibot-22 dataset and 1000 for other datasets. We conduct all experiments on 2 NVIDIA RTX A5000 GPUs.
@@ 5.3 Community-aware Study
To further qualitatively analyze whether the community detection result is meaningful, and how the contrastive learning design helps to deal with the hard samples, we visualize the changes of average cosine similarity among nodes in one community and nodes in different communities. The results shown in Figure 4 prove that the similarities between positive samples continuously rise and those between negative samples decrease, which means same-class nodes are pulled together and diff-class nodes are pushed apart. The results also show that similarities of nodes within and between communities decrease as we dynamically update the community detection module, which proves the community detection can dynamically capture hard samples and the result is helpful for the detection.
Figure 4: Cosine similarity change trends. Positive and negative represent cosine similarities between positive samples and between negative samples, respectively. Within and between represent cosine similarities within one community and between different communities, respectively.
@@ 6 Conclusion
In this paper, we mainly focus on constructing a community-aware graph contrastive learning framework. Despite our best efforts, this paper may still have some limitations. We evaluate our method in three benchmarks, including the most comprehensive dataset Twibot-22 in Twitter bot detection area. However, Hays et al. (2023) still point out that current datasets still have certain limitations in collecting and labeling that affect the evaluation of methods. If we find several communities in real world and record the changes of them during the training process, we can simply reveal the effectiveness of our proposed framework. However, due to our framework is dynamic updated, community structure is dynamic and hard to record.
@@
Our research in this paper follows ethical guidelines and principles. All experiments and data usage are conducted in accordance with the relevant institutional and national guidelines and regulations. Despite the datasets we use contain a large amount of private information existing in social media, we ensure the privacy and confidentiality of any personal or sensitive information used and obtained during the study. Our work aims to contribute positively to online environment, while respecting the dignity, rights, and well-being of individuals and communities. We acknowledge and address any potential biases or ethical considerations in our research.
Acknowledgements
This research is supported by the National Natural Science Foundation of China (No.62106105), the CCF-Baidu Open Fund (No.CCF-Baidu202307), the CCF-Zhipu AI Large Model Fund (No.CCF-Zhipu202315), the Scientific Research Starting Foundation of Nanjing University of Aeronautics and Astronautics (No.YQR21022), the High Performance Computing Platform of Nanjing University of Aeronautics and Astronautics, and the foundation of National College Students’ innovation and entrepreneurship training program (No.202310287085Z).